The Pepper Company plans to invest $500,000 in new equipment and expects to save $95,000 per year. The project is for 8 years and the interest rate is 6%. The equipment has a salvage value of $100,000. The machine hours per year are 2200.

Compute the annual equivalent cash flow per machine hour.
Answer #1

Annual equivalent cash flow = -initial investment (A/P, i%, n) + 95000 + 100000(A/F, i%, n)

= -500000(A/P, 6%, 8) + 95000 + 100000(A/F, 6%, 8)

= -500000*0.16104 + 95000 + 100000*0.10104

= $24586

annual equivalent cash flow per machine hour = $24586/2200 = $11.18 per hour